OpenStudy (anonymous):

-4(x+6)=-40 I got 16 but its wrong ): Can someone show me how to do it Dx I feel dumb

OpenStudy (anonymous):

x + 6 = 10 x = 4

OpenStudy (anonymous):

huh

OpenStudy (anonymous):

-4x -24 = -40 -4x = -16 x = 4

OpenStudy (anonymous):

-4(x+6)=-40 -4x-24=-40 +24 +24 -4x=16 --- --- -4 -4 x=-4 -4(-4+6)=-40 -4(2)=-40 -8=-40 +8 +8 =-32

OpenStudy (anonymous):

oh ok i see what i did wrong thank you (:

OpenStudy (anonymous):

youre welcome